OpenStudy (anonymous):

Calculate the number of proper subsets6,11,10,15

OpenStudy (kmeis002):

A proper subset B of A is a set such that all elements in B are elements in A, but A has atleast one element that is not in B. So if \[A = \left \{ 6, 10, 11, 15 \right \} \] Then \[\left \{ 6 \right \} \] is a proper subset, and \[\left \{ 6,11 \right \} \] \[\left \{ 10,15 \right \} \] Can you list all subsets then?